Problem 05.031 Future Worth Analysis A small strip-mining coal company is trying to decide whether it should purchase or lease a new clamshell. If purchased, the shell will cost $145,000 and is expected to have a $37,500 salvage value after 6 years. Alternatively, the company can lease a clamshell for only $20,000 per year, but the lease payment will have to be made at the beginning of each year. If the clamshell is purchased, it will be leased to other strip-mining companies whenever possible, an activity that is expected to yield revenues of $14,000 per year. If the companys MARR is 18% per year, should the clamshell be purchased or leased on the basis of a future worth analysis? Assume the annual M&O cost is the same for both options. ints eookannual Hint Print The future worth when purchased is $ References The future worth when leased is $ The clamshell should be Corchased-it) .
Answer #1

FW when purchased = -145000(F/P,18%,6) + 14000(F/A,18%,6)+37500

= -145000(2.6995)+14000(9.4419)+37500

= -391427.5+132186.6+37500

= -221740.9

FW when leased = -20000(F/P,18%,6)-20000(F/A,18%,5)(F/P,18%,1)

= -20000(2.6995)-20000(7.1542)(2.6995)

= -53990-386255.25

= -440245.25

The clamshell should be Purchased